OpenStudy (rebecca1588):

The system of equations is coincident.What are the missing values? -2x+5y=-12 -6x+I______I y=I______I

OpenStudy (rebecca1588):

Is this correct -2x+5y=-12, -6x + 15y= -36 ? :)

OpenStudy (rebecca1588):

@superdavesuper am I correct?

OpenStudy (superdavesuper):

very good, @rebecca1588 - u are correct :)

OpenStudy (rebecca1588):

Thanks :)